OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (amistre64):

what is the setup when n=1?

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

2n = n(n+1) , when n=1 then replace n with k in the setup to define when n=k

OpenStudy (amistre64):

the last step will be an induction step using the rule: 2+4+6+....+2n ; when n=k+1 is 2+4+6+....+2(k+1), or better yet 2+4+6+....+2k+2(k+1) Pk + 2(k+1) = k(k+1) + 2(k+1)

OpenStudy (amistre64):

work the right side to match the format of: n(n+1), for n=k+1

OpenStudy (anonymous):

??

OpenStudy (amistre64):

im not real sure what your confusion is .....

OpenStudy (anonymous):

i dont get any of this how do i do for p1?

OpenStudy (amistre64):

replace n with a 1

OpenStudy (anonymous):

then pk2

OpenStudy (anonymous):

pk*

OpenStudy (anonymous):

the pk+1

OpenStudy (amistre64):

P(n) = 2 + 4 + 6 + ... + 2n = n(n+1) P(1) = 2(1) = 1(1+1) ; is true

OpenStudy (amistre64):

for Pk, replace n with a k P(n) = 2 + 4 + 6 + ... + 2n = n(n+1) P(k) = 2 + 4 + 6 + ... + 2k = k(k+1)

OpenStudy (amistre64):

for P(k+1), replace n with a (k+1) ; but we have to show that the right side doesnt change its format P(n) = 2 + 4 + 6 + ... + 2n = n(n+1) P(k+1) = 2 + 4 + 6 + ... + 2k + 2(k+1) ?=? (k+1)((k+1)+1)

OpenStudy (amistre64):

from P(k) we know that: P(k) = 2 + 4 + 6 + ... + 2k = k(k+1) therefore, P(k+1) is just P(k+1) = 2 + 4 + 6 + ... + 2k + 2(k+1) = k(k+1) + 2(k+1) ^^^^^ ^^^^^^ adding 2(k+1) to P(k) P(k+1) = P(k) + 2(k+1) = k(k+1) + 2(k+1) show that : k(k+1) + 2(k+1) = (k+1)((k+1)+1)

OpenStudy (anonymous):

thank you <3

OpenStudy (amistre64):

youre welcome